# FragmentRetro: A Quadratic Retrosynthetic Method Based on Fragmentation Algorithms

### v1

@@ Abstract
Retrosynthesis, the process of deconstructing a target molecule into simpler precursors, is crucial for computer-aided synthesis planning (CASP). Widely adopted tree-search methods often suffer from exponential computational complexity. In this work, we introduce FragmentRetro, a novel retrosynthetic method that leverages fragmentation algorithms, specifically BRICS and r-BRICS, combined with stock-aware exploration and pattern fingerprint screening to achieve quadratic complexity. FragmentRetro recursively combines molecular fragments and verifies their presence in a building block set, providing sets of fragment combinations as retrosynthetic solutions. We present the first formal computational analysis of retrosynthetic methods, showing that tree search exhibits exponential complexity $\mathcal{O}(b^{h})$, DirectMultiStep scales as $\mathcal{O}(h^{6})$, and FragmentRetro achieves $\mathcal{O}(h^{2})$, where $h$ represents the number of heavy atoms in the target molecule and $b$ is the branching factor for tree search. Evaluations on PaRoutes, USPTO-190, and natural products demonstrate that FragmentRetro achieves high solved rates with competitive runtime, including cases where tree search fails. The method benefits from fingerprint screening, which significantly reduces substructure matching complexity. While FragmentRetro focuses on efficiently identifying fragment-based solutions rather than full reaction pathways, its computational advantages and ability to generate strategic starting candidates establish it as a powerful foundational component for scalable and automated synthesis planning.
@@ 1 Introduction
Retrosynthetic analysis, the deconstruction of a target molecule into simpler, typically commercially available precursors, is fundamental to chemical synthesis and drug discovery [1, 2]. Computer-aided synthesis planning (CASP) aims to automate this complex task. A conventional paradigm in CASP involves tree-search algorithms, such as Monte Carlo Tree Search (MCTS) [3], A* search (Retro*) [4], and Depth-First Proof Number (DFPN) search [5]. These methods iteratively apply single-step retrosynthetic (SSR) models, often machine learning-based [6, 7, 8, 9], to explore the vast combinatorial space of possible synthetic routes. While effective, a primary limitation of such tree-search approaches is their inherent exponential computational complexity with respect to tree depth, which can hinder scalability for large and complex targets. Recent advancements aim to mitigate this, for example, by incorporating graph-based neural networks [10], reinforcement learning [11], route context [12], or higher-level abstractions of molecular features  [13]. However, the underlying exponential scaling often persists in the worst-case scenario.
@@
Alternative strategies have emerged to address these scalability challenges. DirectMultiStep (DMS) models, for instance, leverage sequence-to-sequence architectures to generate entire multi-step pathways directly, bypassing iterative search [14]. While DMS has shown promising performance, particularly on benchmarks like PaRoutes [15], and exhibits polynomial complexity, its $\mathcal{O}(h^{6})$ scaling with molecular size $h$ can still be demanding. Other approaches, such as Double-Ended Synthesis Planning (DESP) [16], improve efficiency for specific use cases by conducting a bidirectional tree-search when starting materials are provided by the user, though they still rely on iterative single-step predictions. This landscape motivates the exploration of fundamentally different algorithmic approaches to retrosynthesis that offer improved computational efficiency, particularly for unconstrained initial exploration, without sacrificing solution quality.
@@
In this work, we introduce FragmentRetro, a novel retrosynthetic method that adopts a bottom-up, fragment-based strategy (Fig. 1). FragmentRetro utilizes molecular fragmentation algorithms (BRICS [17] and its revision, r-BRICS [18]) to decompose a target molecule. It then recursively combines these fragments, systematically verifying if the combined fragments exist as substructures within a provided set of commercially available building blocks (BBs). By integrating efficient substructure matching, augmented by pattern fingerprint and property screening, FragmentRetro identifies sets of fragment combinations that reconstruct the target. A key contribution of this paper is the first formal computational complexity analysis comparing these distinct retrosynthetic paradigms. We demonstrate that while tree-search methods scale exponentially ($\mathcal{O}(b^{h})$) and DMS polynomially ($\mathcal{O}(h^{6})$), FragmentRetro achieves a significantly more favorable quadratic complexity ($\mathcal{O}(h^{2})$), albeit with a linear dependency on the stock set size which is effectively managed through parallelization.
@@ 2.1 Tree-search based retrosynthesis
The predominant approach to CASP involves constructing a search tree or graph. Early systems relied on expert-defined chemical rules [1]. Modern methods typically integrate machine learning models for SSR predictions with advanced search algorithms. These SSR models can be template-based, utilizing predefined reaction patterns [6, 19, 7, 8, 20, 21, 22, 9, 23], or template-free, directly predicting reactants from products [24]. Popular search algorithms include Monte Carlo Tree Search (MCTS) [3, 25], A* search Retro* [4], and Proof-Number Search derivatives like DFPN [5].
Recent advancements in tree search focus on improving efficiency and prediction quality. For example, Xie et al. proposed RetroGraph, using graph neural networks to guide search on an explicit graph representation of intermediates to avoid redundant computations [10]. Reinforcement learning techniques, such as in GRASP [11], train agents to learn optimal search policies. FusionRetro [12] aims to improve SSR predictions by incorporating contextual information from the partially built synthetic route. Roh et al. introduced a higher-level retrosynthesis strategy (Higherlev) that abstracts molecular details to simplify the search space, using substructure matching against BBs as a stopping criterion [13]. Despite these innovations, a fundamental challenge for tree-search methods is their worst-case exponential scaling with the size of the target molecule or the depth of the synthetic route, as formally analyzed in this work (Section B).
@@ 2.2 Direct pathway generation and constrained search
Addressing the common real-world scenario where specific starting materials must be utilized, Yu et al. introduced Double-Ended Synthesis Planning (DESP) [16]. DESP employs a bidirectional tree-search approach, simultaneously exploring retrosynthetic steps from the target and forward synthetic steps from user-specified starting materials. This method, guided by a learned synthetic distance cost function, has shown improved efficiency in solving such constrained problems compared to purely unidirectional search. However, like many tree-search methods, DESP combines explicit graph search with learned single-step predictors, and critically, it requires prior human specification of starting materials. Our work, in contrast, aims to identify potential starting material sets automatically as part of its core fragment-based exploration.
@@ 2.3 Molecular fragmentation in chemistry
Molecular fragmentation, the process of breaking molecules into smaller, chemically meaningful pieces, is a well-established technique in cheminformatics. It finds applications in drug design for identifying pharmacophores, in quantitative structure-activity relationship (QSAR) modeling, and as a preprocessing step for various machine learning tasks. The BRICS (Breaking of Retrosynthetically Interesting Chemical Substructures) algorithm [17] is a widely used rule-based method that cleaves molecules at bonds deemed retrosynthetically labile, effectively identifying potential synthons or key disconnections. BRICS has been noted for sometimes producing large, inflexible fragments. To address this, r-BRICS (revised BRICS) was recently developed by Zhang et al., extending the original rule set to enable more granular fragmentation, particularly for challenging structures like fused rings and long aliphatic chains [18]. Our work, FragmentRetro, leverages these fragmentation algorithms not merely for analysis, but as the core engine for a bottom-up retrosynthetic search.
@@ 2.4 Substructure searching
A critical component of FragmentRetro, as well as methods like Higherlev [13], is substructure searching: determining if a given molecular fragment (query) exists within a larger molecule (target, often from a database of BBs). This is a subgraph isomorphism problem. Efficient algorithms like VF2 [26] are commonly used in cheminformatics toolkits such as RDKit [27]. The practical performance of substructure searching can be significantly enhanced by pre-filtering candidates using molecular fingerprints (e.g., pattern fingerprints) or simple property checks (e.g., heavy atom count, ring count), as employed in FragmentRetro. While worst-case complexity for subgraph isomorphism is exponential, these heuristics and optimized algorithms make it feasible for large-scale database searching in practice [28].
@@ 3.1 Fragment-based retrosynthesis
Let $\mathcal{F}_{1}=\{f^{(1)}_{1},f^{(1)}_{2},\ldots,f^{(1)}_{k}\}$ represent the set of $k$ molecular fragments obtained from a fragmentation method (e.g., BRICS [17] and r-BRICS [18]), where each fragment $f^{(1)}_{i}\in\mathcal{F}$ corresponds to a SMILES [29] string. These SMILES strings contain ‘any’ atoms (dummy atoms ‘*’, which represent any atom in SMARTS [30] expressions) each attaching to the fragment bond ends unless no fragmentation is performed on a given SMILES string.
@@
    Initialization: For each fragment $f^{(1)}_{i}\in\mathcal{F}_{1}$, check if $\exists~b\in\mathcal{B}$ s.t. $f^{(1)}_{i}$ is a substructure of $b$, i.e., $f^{(1)}_{i}\subseteq b$. Here, strict substructure matches are done using SMARTS patterns to avoid branchings from non-fragmented sites. Also, the ‘any’ atoms can be hydrogen atoms (no branching from the fragmented sites). We call $f^{(1)}_{i}$ a valid fragment if there is a substructure match. If $\exists~f^{(1)}_{i}\in\mathcal{F}_{1}$ s.t. $f^{(1)}_{i}\not\subseteq b$ $\forall~b\in\mathcal{B}$, the process terminates immediately, as subsequent fragment combinations would also fail to have substructure matches.
@@ 4.1 Computational complexity comparison
The computational complexity analysis is outlined in Section B and summarized in Table 1. The exponential complexity of tree search also applies to the Higherlev approach [13], which employs a higher-level retrosynthetic strategy by abstracting functional groups based on their electronegativity. For the Higherlev method, determining whether a precursor exists in the stock compound set is formulated as a substructure matching task. Since functional group abstraction increases $\Delta h$, their approach effectively reduces $h/\Delta h$ in practice, where $h$ is the number of heavy atoms in the target molecule and $\Delta h$ is the average reduction in heavy atom count per reaction. However, $\Delta h$ remains constant. Consequently, following our derivation in Sections B.1 and B.3, the complexity of their method is given by $\mathcal{O}(C_{fp}\cdot|\mathcal{B}|\cdot b^{h/\Delta h})=\mathcal{O}(b^{h})$ where $b$ is the branching factor of the search tree. For the definition of $C_{fp}$, please refer to Section B.
@@ 4.2 Search performance evaluation
cData collected from the 2.0 version of PaRoutes in their GitHub repository [15] under the Apache-2.0 License.
@@
Table 3: Search performance on USPTO-190 with Buyables as the stock set
Methoda Targetsb,c Solved Rate Run Time (s)d
DMS-Explorer-XL USPTO-190 27.9% 21.9
Original + MCTS [13] USPTO-190 46.3% NA
DMS-Flashe USPTO-190 55.3% 31.3
DMS-Widee USPTO-190 56.8% 105.8
Retro*-0 [4]f USPTO-190 73.2% NA (65.0)
Higherlev + MCTS [13] USPTO-190 73.7% NA
Retro* [4]f USPTO-190 76.8% NA (57.1)
FragmentRetro + BRICSg USPTO-190 53.2% 344.0 (74.2, 49.1, 42.1)
FragmentRetro + r-BRICSg USPTO-190 78.4% 473.1 (102.1, 72.9, 64.5)
a DMS models are run with a beam size of 50 on a single NVIDIA A100 GPU with half-precision floating point inference (FP16). The other methods are run with a single CPU (no parallelization)
b USPTO-190 is from [4] (downloaded from [31] under the CC BY 4.0 License).
c The stock compounds (Buyables) are from [13] under the CC BY 4.0 License and includes 0.329M buyable building blocks from eMolecules, Sigma-Aldrich, Mcule, ChemBridge Hit2Lead, and WuXi LabNetwork.
@@
f Evaluated in this work using Retro*’s [4] official repository. The original checkpoint is used with a maximum of 500 iterations. Runtimes in parentheses are first-solution times, as the repository lacks search continuation.
@@
Table 2 presents the search performance of various methods on the PaRoutes test sets n1 and n5. All methods were evaluated using a single CPU/core (no parallelization), except for the DMS models, which were run on an NVIDIA A100 GPU. The stock set for each test follows the SMs defined in n1 or n5, as done in [15]. Among the evaluated methods, Retro*, a tree-search approach, achieved the highest solved rate. The DMS Explorer-XL model [14] exhibited competitive performance, solving a significant portion of the test set while providing a 20x speedup compared to tree-search methods. Similarly, FragmentRetro demonstrated decent performance, particularly with r-BRICS fragmentation, which outperformed DMS in solved rate while achieving a 30x speedup over tree-search methods. Table 2 also reports the median number of solutions. Since the PaRoutes paper [15] clustered synthetic routes using the method in [32] to estimate route diversity, we compare the number of solutions from FragmentRetro to the number of clusters from PaRoutes. However, these metrics are not directly comparable, as a single FragmentRetro solution can correspond to multiple synthetic routes—each valid fragment combination may have substructure matches with different BBs.
@@
Table 3 summarizes the search performance of different methods on the USPTO-190 dataset, using the Buyables stock set (0.329M BBs) provided by the Higherlev study [13]. This stock set is a more practical choice than the eMolecule screening compounds (23.1M compounds) used in Retro* [4], as many screening compounds require custom synthesis. It is important to note that 13 compounds from USPTO-190 appear as target compounds in the single-step training dataset for Higherlev, while 47 compounds appear as targets or intermediates in the PaRoutes training set used by the DMS models. The DMS study reports solved rates both with and without these 47 compounds, showing minor differences. We expect similar results for Higherlev. FragmentRetro with BRICS fragmentation achieved a comparable solved rate to the best-performing DMS models but did not surpass Higherlev. However, with r-BRICS fragmentation, FragmentRetro achieved the highest solved rate among all methods. Despite this, the runtime of FragmentRetro is similar to that of tree-search methods in Table 2, as execution time scales with the stock set size ($|\mathcal{B}|$), even with the prefactor $C_{fp}$. This is due to substructure matching being the dominant computational bottleneck, as each candidate BB must be checked against the fragments. Fortunately, these checks are independent and can be efficiently parallelized across multiple CPU cores. To demonstrate this scalability, we also report FragmentRetro runtime using 5, 10, and 20 CPU cores. The results show that parallelization significantly reduces runtime, achieving speeds comparable to or better than DMS models. Notably, 5 and 10 CPU cores exhibit near-perfect parallelization efficiency, with speedups of 4.6x and 6.9x, respectively. However, with 20 CPU cores, the speedup is only 8.2x, indicating diminishing returns. This suggests that at this stock set size, the tradeoff between parallelization efficiency and overhead is optimal around 10 CPU cores.

### v2

@@ Abstract
Retrosynthesis, the process of deconstructing a target molecule into simpler precursors, is crucial for computer-aided synthesis planning (CASP). Widely adopted tree-search methods often suffer from exponential computational complexity. In this work, we introduce FragmentRetro, a novel retrosynthetic method that leverages fragmentation algorithms, specifically BRICS and r-BRICS, combined with stock-aware exploration and pattern fingerprint screening to achieve quadratic complexity. FragmentRetro recursively combines molecular fragments and verifies their presence in a building block set, providing sets of fragment combinations as retrosynthetic solutions. We present the first formal computational analysis of retrosynthetic methods, showing that tree search exhibits exponential complexity $\mathcal{O}(b^{h})$, DirectMultiStep scales as $\mathcal{O}(h^{6})$, and FragmentRetro achieves $\mathcal{O}(h^{2})$, where $h$ represents the number of heavy atoms in the target molecule and $b$ is the branching factor for tree search. Evaluations on PaRoutes, USPTO-190, and natural products demonstrate that FragmentRetro achieves high solved rates with competitive runtime, including cases where tree search fails. The method benefits from fingerprint screening, which significantly reduces substructure matching complexity. While FragmentRetro focuses on efficiently identifying fragment-based solutions rather than full reaction pathways, its computational advantages and ability to generate strategic starting candidates establish it as a powerful foundational component for scalable and automated synthesis planning. FragmentRetro code is available on GitHub
@@ 1 Introduction
Retrosynthetic analysis, the deconstruction of a target molecule into simpler, typically commercially available precursors, is fundamental to chemical synthesis and drug discovery [casp_1969, grzy_2016]. Computer-aided synthesis planning (CASP) aims to automate this complex task. A conventional paradigm in CASP involves tree-search algorithms, such as Monte Carlo Tree Search (MCTS) [mcts_2018], A* search (Retro*) [retrostar], and Depth-First Proof Number (DFPN) search [dfpn]. These methods iteratively apply single-step retrosynthetic (SSR) models, often machine learning-based [coley_aug_2020, coley_retrosim_2017, ishida_gcn_2019, site_spec_2024], to explore the vast combinatorial space of possible synthetic routes. While effective, a primary limitation of such tree-search approaches is their inherent exponential computational complexity with respect to tree depth, which can hinder scalability for large and complex targets. Recent advancements aim to mitigate this, for example, by incorporating graph-based neural networks [retrograph], reinforcement learning [grasp_2022], route context [fusionretro_2023], or higher-level abstractions of molecular features  [higherlev_2025]. However, the underlying exponential scaling often persists in the worst-case scenario.
@@
Alternative strategies have emerged to address these scalability challenges. DirectMultiStep (DMS) models, for instance, leverage sequence-to-sequence architectures to generate entire multi-step pathways directly, bypassing iterative search [dms_2025]. While DMS has shown promising performance, particularly on benchmarks like PaRoutes [paroutes], and exhibits polynomial complexity, its $\mathcal{O}(h^{6})$ scaling with molecular size $h$ can still be demanding. Other approaches, such as Double-Ended Synthesis Planning (DESP) [desp_2024], improve efficiency for specific use cases by conducting a bidirectional tree-search when starting materials are provided by the user, though they still rely on iterative single-step predictions. This landscape motivates the exploration of fundamentally different algorithmic approaches to retrosynthesis that offer improved computational efficiency, particularly for unconstrained initial exploration, without sacrificing solution quality.
@@
In this work, we introduce FragmentRetro, a novel retrosynthetic method that adopts a bottom-up, fragment-based strategy (Fig. 1). FragmentRetro utilizes molecular fragmentation algorithms (BRICS [brics_2008] and its revision, r-BRICS [r_brics_2024]) to decompose a target molecule. It then recursively combines these fragments, systematically verifying if the combined fragments exist as substructures within a provided set of commercially available building blocks (BBs). By integrating efficient substructure matching, augmented by pattern fingerprint and property screening, FragmentRetro identifies sets of fragment combinations that reconstruct the target. A key contribution of this paper is the first formal computational complexity analysis comparing these distinct retrosynthetic paradigms. We demonstrate that while tree-search methods scale exponentially ($\mathcal{O}(b^{h})$) and DMS polynomially ($\mathcal{O}(h^{6})$), FragmentRetro achieves a significantly more favorable quadratic complexity ($\mathcal{O}(h^{2})$), albeit with a linear dependency on the stock set size which is effectively managed through parallelization.
@@ 2.1 Tree-search based retrosynthesis
The predominant approach to CASP involves constructing a search tree or graph. Early systems relied on expert-defined chemical rules [casp_1969]. Modern methods typically integrate machine learning models for SSR predictions with advanced search algorithms. These SSR models can be template-based, utilizing predefined reaction patterns [coley_aug_2020, coley_gcn_2019, coley_retrosim_2017, ishida_gcn_2019, hopfield_ssr, segler_ml_2017, localretro_2021, site_spec_2024, coley_rank_2017], or template-free, directly predicting reactants from products [wlnet_coley_2017]. Popular search algorithms include Monte Carlo Tree Search (MCTS) [mcts_2018, aizynthfinder_2020], A* search Retro* [retrostar], and Proof-Number Search derivatives like DFPN [dfpn].
Recent advancements in tree search focus on improving efficiency and prediction quality. For example, retrograph proposed RetroGraph, using graph neural networks to guide search on an explicit graph representation of intermediates to avoid redundant computations [retrograph]. Reinforcement learning techniques, such as in GRASP [grasp_2022], train agents to learn optimal search policies. FusionRetro [fusionretro_2023] aims to improve SSR predictions by incorporating contextual information from the partially built synthetic route. higherlev_2025 introduced a higher-level retrosynthesis strategy (Higherlev) that abstracts molecular details to simplify the search space, using substructure matching against BBs as a stopping criterion [higherlev_2025]. Despite these innovations, a fundamental challenge for tree-search methods is their worst-case exponential scaling with the size of the target molecule or the depth of the synthetic route, as formally analyzed in this work (Section B).
@@ 2.2 Direct pathway generation and constrained search
Addressing the common real-world scenario where specific starting materials must be utilized, desp_2024 introduced Double-Ended Synthesis Planning (DESP) [desp_2024]. DESP employs a bidirectional tree-search approach, simultaneously exploring retrosynthetic steps from the target and forward synthetic steps from user-specified starting materials. This method, guided by a learned synthetic distance cost function, has shown improved efficiency in solving such constrained problems compared to purely unidirectional search. However, like many tree-search methods, DESP combines explicit graph search with learned single-step predictors, and critically, it requires prior human specification of starting materials. Our work, in contrast, aims to identify potential starting material sets automatically as part of its core fragment-based exploration.
@@ 2.3 Molecular fragmentation in chemistry
Molecular fragmentation, the process of breaking molecules into smaller, chemically meaningful pieces, is a well-established technique in cheminformatics. It finds applications in drug design for identifying pharmacophores, in quantitative structure-activity relationship (QSAR) modeling, and as a preprocessing step for various machine learning tasks. The BRICS (Breaking of Retrosynthetically Interesting Chemical Substructures) algorithm [brics_2008] is a widely used rule-based method that cleaves molecules at bonds deemed retrosynthetically labile, effectively identifying potential synthons or key disconnections. BRICS has been noted for sometimes producing large, inflexible fragments. To address this, r-BRICS (revised BRICS) was recently developed by r_brics_2024, extending the original rule set to enable more granular fragmentation, particularly for challenging structures like fused rings and long aliphatic chains [r_brics_2024]. Our work, FragmentRetro, leverages these fragmentation algorithms not merely for analysis, but as the core engine for a bottom-up retrosynthetic search.
@@ 2.4 Substructure searching
A critical component of FragmentRetro, as well as methods like Higherlev [higherlev_2025], is substructure searching: determining if a given molecular fragment (query) exists within a larger molecule (often from a database of BBs). This is a subgraph isomorphism problem. Efficient algorithms like VF2 [substruct_algo_2004] are commonly used in cheminformatics toolkits such as RDKit [rdkit_2006]. The practical performance of substructure searching can be significantly enhanced by pre-filtering candidates using molecular fingerprints (e.g., pattern fingerprints) or simple property checks (e.g., heavy atom count, ring count), as employed in FragmentRetro. While worst-case complexity for subgraph isomorphism is exponential, these heuristics and optimized algorithms make it feasible for large-scale database searching in practice [substruct_bench_2012].
@@ 3.1 Fragment-based retrosynthesis
Let $\mathcal{F}_{1}=\{f^{(1)}_{1},f^{(1)}_{2},\ldots,f^{(1)}_{k}\}$ represent the set of $k$ molecular fragments obtained from a fragmentation method (e.g., BRICS [brics_2008] and r-BRICS [r_brics_2024]), where each fragment $f^{(1)}_{i}\in\mathcal{F}$ corresponds to a SMILES [smiles] string. These SMILES strings contain ‘any’ atoms (dummy atoms ‘*’, which represent any atom in SMARTS [smarts] expressions) each attaching to the fragment bond ends unless no fragmentation is performed on a given SMILES string.
@@
    Initialization: For each fragment $f^{(1)}_{i}\in\mathcal{F}_{1}$, check if $\exists~b\in\mathcal{B}$ s.t. $f^{(1)}_{i}$ is a substructure of $b$, i.e., $f^{(1)}_{i}\subseteq b$. Here, strict substructure matches are done using SMARTS patterns to avoid branchings from non-fragmented sites. Also, the ‘any’ atoms can be hydrogen atoms (no branching from the fragmented sites). We call $f^{(1)}_{i}$ a valid fragment if there is a substructure match. Aromaticity and chirality must match for a candidate BB to be considered as containing the substructure of the fragment, except at the fragmented sites where stereochemistry may be undefined. If $\exists~f^{(1)}_{i}\in\mathcal{F}_{1}$ s.t. $f^{(1)}_{i}\not\subseteq b$ $\forall~b\in\mathcal{B}$, the process terminates immediately, as subsequent fragment combinations would also fail to have substructure matches.
@@ 3.3 Computational complexity comparison
The computational complexity analysis is outlined in Section B and summarized in Table 1. Here, we use the heavy atom count of a given target $h$ as a proxy for problem difficulty. For tree-search methods, the depth of the search tree for a given target can be estimated as $h/\Delta h$, where $\Delta h$ is the average reduction in heavy-atom count per reaction. Therefore, tree-search methods scale exponentially as $\mathcal{O}(b^{h/\Delta h})=\mathcal{O}(b^{h})$, where $b$ is the branching factor of the search tree and $\Delta h$ can be treated as constant. In comparison, DirectMultiStep models require repeated computation of attention matrices and exhibit polynomial complexity of $\mathcal{O}(h^{6})$, since the route sequence length scales as $\mathcal{O}(h^{2})$ and each attention operation scales as $\mathcal{O}(h^{4})$.
For FragmentRetro, the dominant computational cost arises from substructure matching. However, fingerprint screening eliminates the majority of non-matching candidates before the substructure check, effectively keeping this process inexpensive in practice. This introduces a constant prefactor $C_{fp}$ representing the filtering efficiency. The cost still scales linearly with the stock set size $|\mathcal{B}|$, and the number of possible fragment combinations scales quadratically with $h$. Therefore, the overall complexity of FragmentRetro can be derived as $\mathcal{O}(C_{fp}\cdot|\mathcal{B}|\cdot h^{2})=\mathcal{O}(h^{2})$. Note that this analysis describes the asymptotic behavior and comes with certain caveats. For example, we assume an ideal fingerprint that prevents collisions between molecular features, but in practice the fingerprint size is finite. Additionally, the worst case could occur for extremely ring-heavy target molecules, where r-BRICS fragmentation of ring bonds produces a large number of fragments and combinations, increasing the computational cost.
@@ 4.1 Search performance evaluation
cData collected from the 2.0 version of PaRoutes in their GitHub repository [paroutes] under the Apache-2.0 License.
@@
Table 3: Search performance on USPTO-190a with Buyablesb as the stock set
Methodc Solved Rate Run Time (s)d
DMS-Explorer-XL 27.9% 21.9
Original + MCTS [higherlev_2025] 46.3% NA
DMS-Flashe 55.3% 31.3
DMS-Widee 56.8% 105.8
Retro*-0 [retrostar]f 73.2% NA (65.0)
Higherlev + MCTS [higherlev_2025] 73.7% NA
Retro* [retrostar]f 76.8% NA (57.1)
FragmentRetro + BRICSg 53.2% 344.0 (74.2, 49.1, 42.1)
FragmentRetro + r-BRICSg 78.4% 473.1 (102.1, 72.9, 64.5)
a USPTO-190 is from [retrostar] (downloaded from [shee2025figshare] under the CC BY 4.0 License).
b The stock compounds (Buyables) are from [higherlev_2025] under the CC BY 4.0 License and includes 0.329M buyable building blocks from eMolecules, Sigma-Aldrich, Mcule, ChemBridge Hit2Lead, and WuXi LabNetwork.
c DMS models are run with a beam size of 50 on a single NVIDIA A100 GPU with half-precision floating point inference (FP16). The other methods are run with a single CPU (no parallelization)
@@
f Evaluated in this work using Retro*’s [retrostar] official repository. The original checkpoint is used with a maximum of 500 iterations. Runtimes in parentheses are first-solution times, as the repository lacks search continuation.
@@
Table 2 presents the search performance of various methods on the PaRoutes test sets n1 and n5. All methods were evaluated using a single CPU/core (no parallelization), except for the DMS models, which were run on an NVIDIA A100 GPU. The stock set for each test follows the SMs defined in n1 or n5, as done in [paroutes]. Among the evaluated methods, Retro*, a tree-search approach, achieved the highest solved rate. The DMS Explorer-XL model [dms_2025] exhibited competitive performance, solving a significant portion of the test set while providing a 20x speedup compared to tree-search methods. Similarly, FragmentRetro demonstrated decent performance, particularly with r-BRICS fragmentation, which outperformed DMS in solved rate while achieving a 30x speedup over tree-search methods. For FragmentRetro, a target is considered solved if at least one solution is found, where the solution corresponds to a set of valid fragments that can reconstruct the target compound, since the method does not output DAGs. Table 2 also reports the median number of solutions. Since the PaRoutes paper [paroutes] clustered synthetic routes using the method in [route_dist_2022] to estimate route diversity, we compare the number of solutions from FragmentRetro to the number of clusters from PaRoutes. However, these metrics are not directly comparable, as a single FragmentRetro solution can correspond to multiple synthetic routes—each valid fragment combination may have substructure matches with different BBs.
@@
Table 3 summarizes the search performance of different methods on the USPTO-190 dataset, using the Buyables stock set (0.329M BBs) provided by the Higherlev study [higherlev_2025]. This stock set is a more practical choice than the eMolecule screening compounds (23.1M compounds) used in Retro* [retrostar], as many screening compounds require custom synthesis. It is important to note that 13 compounds from USPTO-190 appear as target compounds in the single-step training dataset for Higherlev, while 47 compounds appear as targets or intermediates in the PaRoutes training set used by the DMS models. The DMS study reports solved rates both with and without these 47 compounds, showing minor differences. We expect similar results for Higherlev. FragmentRetro with BRICS fragmentation achieved a comparable solved rate to the best-performing DMS models but did not surpass Higherlev. However, with r-BRICS fragmentation, FragmentRetro achieved the highest solved rate among all methods. Despite this, the runtime of FragmentRetro is similar to that of tree-search methods in Table 2, as execution time scales with the stock set size ($|\mathcal{B}|$), even with the prefactor $C_{fp}$. This is due to substructure matching being the dominant computational bottleneck, as each candidate BB must be checked against the fragments. Fortunately, these checks are independent and can be efficiently parallelized across multiple CPU cores. To demonstrate this scalability, we also report FragmentRetro runtime using 5, 10, and 20 CPU cores. The results show that parallelization significantly reduces runtime, achieving speeds comparable to or better than DMS models. Notably, 5 and 10 CPU cores exhibit near-perfect parallelization efficiency, with speedups of 4.6x and 6.9x, respectively. However, with 20 CPU cores, the speedup is only 8.2x, indicating diminishing returns. This suggests that at this stock set size, the tradeoff between parallelization efficiency and overhead is optimal around 10 CPU cores.